# Pushing the limits of unconstrained machine-learned interatomic potentials

### v1

@@ I Introduction
The past decade has witnessed the widespread adoption of machine-learned interatomic potentials (MLIPs) [3]. By training on reference data from quantum-mechanical calculations, MLIPs achieve accuracy approaching first-principles methods at highly reduced computational cost. While initial developments focused on system-specific training, the availability of increasingly diverse datasets [4] – spanning much of the periodic table and containing millions of labeled configurations – has driven the training of general-purpose (or universal) MLIPs [5], which can afford good-quality predictions across very diverse systems. This shift has encouraged the adoption of more expressive architectures, notably graph neural networks [6], as opposed to earlier physically inspired functional forms which might be too rigid or limited [7]. It has even prompted some models to discard explicit physical symmetries in favor of unconstrained architectures where these are learned from the training data [8, 9]. While unconstrained architectures can achieve high computational efficiency, most often by not enforcing rotational symmetries and/or conservation of energy, models trained for practitioners and popular benchmarks [10] often rely on explicitly enforcing all physical symmetries of the learning target.
@@ II.3 Applications and benchmarks
Even though the ultimate test of the utility of a MLIP is whether it can be used reliably for practical atomistic simulation tasks, it is often useful to have standardized benchmarks to encourage the development of more accurate and/or efficient MLIPs. Matbench-discovery [10] was among the first to be introduced, and it features the largest number of state-of-the-art architectures, but it does not contain tests on molecular dynamics applications. Very recently, LAMbench [36] and MLIP Arena [37] have also been proposed. These benchmark suites focus almost exclusively on materials, and no analogues exist for molecules, to the best of our knowledge and at the time of writing. As a result, some practitioners evaluate model accuracy on the various test splits of the SPICE-MACE-OFF dataset [26, 38] to benchmark models for applications in organic chemistry and biochemistry.
@@ III Theory and methods
The question of whether a physical prior should be built into the functional form of a MLIP or learned as part of the training process is not one with a simple answer. The best strategy depends on how constraining the functional form affects the expressivity of the architecture, as well as its computational cost – which in turn depends on how well the necessary mathematical operations are supported by modern libraries and hardware. It also important to consider how hard it is to learn the symmetry, and to monitor and correct for it at inference time. Looking at the current landscape, modern GNN-based MLIPs almost universally enforce permutation and translation symmetries, while several recent models relax the requirement for rotational symmetry and energy conservation.
@@ IV.1 Large-scale materials datasets
We apply the proposed architecture to three large-scale materials databases – the MPtrj dataset [4], the subsampled Alexandria dataset [21, 22] and OMat24 [22] – to demonstrate its behavior in the large-data regime and to compare it to state-of-the-art equivariant MLIPs. This “PET-OAM” model (trained on all three datasets) contains 730M parameters. Thanks to the increased size of the node representations and the unconstrained architecture, this large parameter counts can be achieved without sacrificing computational efficiency (see inference timings in Sec. IV.2).
The model were pre-trained using non-conservative forces and stresses, then fine-tuned using their conservative counterparts. This two-step procedure was shown to save computational time in Refs. [31] and [30], often also resulting in better accuracies. Compared to other published training protocols [19, 16, 30], we find that our rotationally unconstrained architecture takes more passes through the data to train to convergence (although this does not necessarily translate to a higher computational cost, see Sec. IV.2). This is not surprising: as argued in Sec. III, an unconstrained model needs to learn the equivariance constraints during training. In contrast, trained unconstrained models can be fine-tuned quickly: for example, as shown in App. B, a model trained on OMat24 can be fine-tuned to the smaller MAD dataset in fewer than 1/20th of the epochs needed to train from scratch, while achieving 50% lower test error than training on MAD alone.
@@ IV.2 The SPICE molecular benchmark
Unlike the material domain, applications of MLIPs to molecular systems have not yet seen the development of established benchmark suites. One of the most widely used large-scale MLIP benchmark in the molecular domain is the test accuracy on the SPICE [26] dataset, in the version proposed by Ref. [38].
IV.2.1 Training and accuracy
@@ V Discussion
Overall, this work shows that the PET architecture, which had already proven itself as a lightweight model trained on a comparatively small dataset [18], retains excellent performance at scale. It establishes rotationally unconstrained architectures as promising alternatives to achieve better trade-offs between accuracy, parameter budget, and computational cost in general-purpose MLIPs and their applications. The pre-trained models we make available with this work span a broad range of parameter counts and computational cost, and can be used by practitioners to quickly fine-tune models on bespoke datasets, facilitating applications, as well as testing the PET architecture on more challenging use cases beyond the benchmarks we present. Given that, in this domain, unconstrained models has received much less attention than equivariant neural networks, we are convinced that there is substantial potential for further improvements in both accuracy and efficiency.
@@ VI Data availability statement
The data that support the findings of this study are openly available. The PET models are freely available at https://github.com/lab-cosmo/upet/, and can be run and fine-tuned using tools available from http://metatensor.org/. These are also documented in Ref. 55.
@@ Acknowledgements.
The authors would like to thank M. Langer, J. Schmidt, B. Wood and M. Gao for helpful discussions. AM and MC acknowledge support from an Industrial Grant from BASF and from EPFL. FB was supported by a project within the Platform for Advanced Scientific Computing (PASC), and by the MARVEL National Centre of Competence in Research (NCCR), funded by the Swiss National Science Foundation (SNSF, grant number 182892) PP and MC acknowledge the funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101001890-FIAMMA). This work was supported by the Swiss AI Initiative (2025 Fellowship Program).
@@ A.3 Training
The official training/validation/test splits were used for the OMat24 [22], sAlex [22], SPICE [38] and MAD [27] datasets. Since the MPtrj dataset [4] does not have official splits, as far as we are aware, we added it in its entirety to the training set for the OAM fine-tuning run.
@@ A.4 Hyperparameters
The hyperparameters employed to train the models presented in this work are shown in Tab. 3.
Table 3: Model and training hyperparameters for the models evaluated in this work. A linear learning-rate warm-up of 10% of the total number of epochs was performed in all cases. The number of node features corresponds to four times that of edge features for all models.
@@ Appendix C Matbench-discovery evaluation
Tab. 6 reports the performance of different models on the phonons task of matbench-discovery, which evaluates the symmetric mean relative error on lattice thermal conductivity [57], for varying finite-difference displacements in force-constant calculations. All results use central finite differences. We find that this choice is crucial for non-equivariant models (and it is also adopted for ORB models), as it can more than halve the error by averaging out small inconsistencies between symmetry-equivalent displacements (see also App. G). Our results show little variability with the displacement size, with optimal values slightly larger than the default 0.03 Å used in phonopy (and matbench-discovery). A similar investigation in Ref. [30] also found that the 0.03 Å standard is not optimal in general. Overall, the conservative model achieves state-of-the-art performance, while the non-conservative model remains competitive with conservative approaches and outperforms other direct-force models on the leaderboard.
@@ Appendix D LAMbench evaluation
The conservative PET-OAM model achieves the best overall accuracy in predicting the shear modulus (Table 10), and competes closely with other universal models in predicting the bulk modulus. Upon comparing the results of the conservative (PET-OAM-C and PET-MPtrj-C) models against their non-conservative analogs (PET-OAM-NC) and (PET-MPtrj-NC), we clearly see that elastic moduli are very sensitive to the lack of energy conservation. This comes as no surprise, as geometry optimization is part of the protocol for calculating elastic moduli, and this it can fail upon having no relaxation constraints and no rotational averaging, as we show in Sections IV.1 and G. MAD-trained models (PET-MAD and PET-OMAD) show comparatively worse results due to a difference in the underlying DFT. This observation also indicates that some of the materials properties (like the elastic moduli) can be actually very sensitive to a choice of DFT flavor and settings.
@@ E F E F E F E F E F E F
Appendix G Details on geometry optimization and phonon calculations

### v2

@@ I Introduction
The past decade has witnessed the widespread adoption of machine-learned interatomic potentials (MLIPs) [3]. By training on reference data from quantum-mechanical calculations, MLIPs achieve accuracy approaching first-principles methods at highly reduced computational cost. While initial developments focused on system-specific training, the availability of increasingly diverse datasets [4] – spanning much of the periodic table and containing millions of labeled configurations – has driven the training of general-purpose (or universal) MLIPs [5], which can afford good-quality predictions across very diverse systems. This shift has encouraged the adoption of more expressive architectures, notably graph neural networks [6], as opposed to earlier physically inspired functional forms which might be too rigid or limited [7]. It has even prompted some models to discard explicit physical symmetries in favor of unconstrained architectures where these are learned from the training data [8, 9, 10]. While unconstrained architectures can achieve high computational efficiency, most often by not enforcing rotational symmetries and/or conservation of energy, models trained for practitioners and popular benchmarks [11] often rely on explicitly enforcing all physical symmetries of the learning target.
@@ II.3 Applications and benchmarks
Even though the ultimate test of the utility of a MLIP is whether it can be used reliably for practical atomistic simulation tasks, it is often useful to have standardized benchmarks to encourage the development of more accurate and/or efficient MLIPs. Matbench-discovery [11] was among the first to be introduced, and it features the largest number of state-of-the-art architectures, but it does not contain tests on molecular dynamics applications. Very recently, LAMbench [37] and MLIP Arena [38] have also been proposed. In order to benchmark models for applications in organic chemistry and biochemistry, some practitioners evaluate model accuracy on the various test splits of the SPICE-MACE-OFF dataset [27, 39], and the Open Molecules leaderboard and community challenges [24] are also increasingly used in this domain.
@@ III Theory and methods
The question of whether a physical prior should be built into the functional form of a MLIP or learned as part of the training process is not one with a simple answer. The best strategy depends on how constraining the functional form affects the expressivity of the architecture, as well as its computational cost – which in turn depends on how well the necessary mathematical operations are supported by modern libraries and hardware. It is also important to consider how hard it is to learn the symmetry, and to monitor and correct for it at inference time. Looking at the current landscape, modern GNN-based MLIPs almost universally enforce permutation and translation symmetries, while several recent models relax the requirement for rotational symmetry and energy conservation.
@@ IV.1 Large-scale materials datasets
We apply the proposed architecture to three large-scale materials databases – the MPtrj dataset [4], the subsampled Alexandria dataset [22, 23] and OMat24 [23] – to demonstrate its behavior in the large-data regime and to compare it to state-of-the-art equivariant MLIPs. This “PET-OAM” model (trained on all three datasets) contains 730M parameters. Thanks to the increased size of the node representations and the unconstrained architecture, this large parameter count can be achieved without sacrificing computational efficiency (see inference timings in Sec. IV.2).
The model was pre-trained using non-conservative forces and stresses, then fine-tuned using their conservative counterparts. This two-step procedure was shown to save computational time in Refs. [32] and [31], often also resulting in better accuracies. Compared to other published training protocols [20, 17, 31], we find that our rotationally unconstrained architecture takes more passes through the data to train to convergence (although this does not necessarily translate to a higher computational cost, see Sec. IV.2). This is not surprising: as argued in Sec. III, an unconstrained model needs to learn the equivariance constraints during training. In contrast, trained unconstrained models can be fine-tuned quickly: for example, as shown in App. B, a model trained on OMat24 can be fine-tuned to the smaller Massive Atomistic Diversity (MAD) dataset  [53] in fewer than 1/20th of the epochs needed to train from scratch  [19] , while achieving 50% lower test error than training on MAD alone.
@@ IV.2 The SPICE molecular benchmark
Unlike the material domain, applications of MLIPs to molecular systems have not yet seen the development of established benchmark suites. Two of the most widely used large-scale MLIP benchmarks in the molecular domain are the test accuracy on the SPICE [27] dataset (in the version proposed by Ref. [39]) and the Open Molecules leaderboard and community challenges [24]. In this section, we focus on the former; training and evaluation exercises on the large-scale OMol-1 dataset [24] are available in App. G.
Training and accuracy
@@ V Discussion
Overall, this work shows that the PET architecture, which had already proven itself as a lightweight model trained on a comparatively small dataset [19], retains excellent performance at scale. It establishes rotationally unconstrained architectures as promising alternatives to achieve better trade-offs between accuracy, parameter budget, and computational cost in general-purpose MLIPs and their applications. The pre-trained models we make available with this work span a broad range of parameter counts and computational cost, and can be used by practitioners to quickly fine-tune models on bespoke datasets, facilitating applications, as well as testing the PET architecture on more challenging use cases beyond the benchmarks we present. Given that, in this domain, unconstrained models have received much less attention than equivariant neural networks, we are convinced that there is substantial potential for further improvements in both accuracy and efficiency.
@@ VI Data availability statement
The data that support the findings of this study are openly available. The PET models are freely available at https://github.com/lab-cosmo/upet/, and they can be run and fine-tuned using tools available from https://metatensor.org/, which are also documented in Ref. 57.
Examples of how to run geometry optimization and phonon calculations with unconstrained models can be found at https://atomistic-cookbook.org/examples/pet-relaxation/pet-relaxation.html and https://atomistic-cookbook.org/examples/pet-phonons/pet-phonons.html, respectively.
@@ Acknowledgements.
The authors would like to thank M. Langer, B. Wood and M. Gao for helpful discussions. AM and MC acknowledge support from an Industrial Grant from BASF and from EPFL. FB was supported by a project within the Platform for Advanced Scientific Computing (PASC), and by the MARVEL National Centre of Competence in Research (NCCR), funded by the Swiss National Science Foundation (SNSF, grant number 182892) PP and MC acknowledge the funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101001890-FIAMMA). This work was supported by the Swiss AI Initiative (2025 Fellowship Program). JS acknowledges support from the Swiss National Science Foundation (SNSF) under Ambizione grant number 233444. Computing resources were also provided by the Swiss National Supercomputing Centre under project no. lp26, lp95 and lp133.
@@ A.1 Architecture
Finally, we would like to stress that all experiments presented in this manuscript do not make use of the Equivariant Coordinate System Ensemble (ECSE) [9] to enforce exact energy invariance.
@@ A.3 Training
The official training/validation/test splits were used for the OMat24 [23], sAlex [23], SPICE [39], MAD [28] and OMol25 [24] datasets. Since the MPtrj dataset [4] does not have official splits, as far as we are aware, we added it in its entirety to the training set for the OAM fine-tuning run.
See Sec. G for model and training details relative to the OMol models.
@@ A.4 Hyperparameters
The hyperparameters employed to train the models presented in this work are shown in Tab. 3. See Sec. G for model and training hyperparameters relative to the OMol models.
Table 3: Model and training hyperparameters for the models evaluated in this work. A linear learning-rate warm-up of 10% of the total number of epochs was performed in all cases. The number of node features corresponds to four times that of edge features for all models. See Sec. G for model and training hyperparameters relative to the OMol models.
@@ Appendix C Matbench-discovery evaluation
Tab. 6 reports the performance of different models on the phonons task of matbench-discovery, which evaluates the symmetric mean relative error on lattice thermal conductivity [59], for varying finite-difference displacements in force-constant calculations. All results use central finite differences. We find that this choice is crucial for non-equivariant models (and it is also adopted for ORB models), as it can more than halve the error by averaging out small inconsistencies between symmetry-equivalent displacements (see also App. H). Our results show little variability with the displacement size, with optimal values slightly larger than the default 0.03 Å used in phonopy (and matbench-discovery). A similar investigation in Ref. [31] also found that the 0.03 Å standard is not optimal in general. Overall, the conservative model achieves state-of-the-art performance, while the non-conservative model remains competitive with conservative approaches and outperforms other direct-force models on the leaderboard.
@@ Appendix D LAMbench evaluation
The conservative PET-OAM model achieves the best overall accuracy in predicting the shear modulus (Table 10), and competes closely with other universal models in predicting the bulk modulus. Upon comparing the results of the conservative (PET-OAM-C and PET-MPtrj-C) models against their non-conservative analogs (PET-OAM-NC) and (PET-MPtrj-NC), we clearly see that elastic moduli are very sensitive to the lack of energy conservation. This comes as no surprise, as geometry optimization is part of the protocol for calculating elastic moduli, and this it can fail upon having no relaxation constraints and no rotational averaging, as we show in Sections IV.1 and H. MAD-trained models (PET-MAD and PET-OMAD) show comparatively worse results due to a difference in the underlying DFT. This observation also indicates that some of the materials properties (like the elastic moduli) can be actually very sensitive to a choice of DFT flavor and settings.
@@ E F E F E F E F E F E F
Appendix G The Open Molecules Benchmark
We trained additional models on the OMol-1 dataset [24] containing 140M systems. To accommodate the varying charge states and spin multiplicities present in the dataset, we introduced a system-conditioning embedding into the PET architecture that closely follows the embeddings introduced in Ref. [24]. Separate embeddings for charge and spin states, each matching the dimensionality of the node features, are projected onto the node dimensions via an additional linear layer. These embeddings are then added to the existing node embeddings at each message-passing step. We initialize the projection layer to zero to stabilize training and to enable seamless combination of the system conditioning with already-pretrained models. During training, energies were pre-normalized by subtracting atomic energies using the neutral atomic reference values provided with the dataset, followed by additional linear fit. To enable efficient training despite the varying molecule sizes in OMol, we use max-atom batching rather than a fixed system number. Similar to the implementation in the fairchem package, we greedily fill batches until the next system would exceed the maximum number of atoms per batch, then start a new batch. Batches are created once, and their order is shuffled each epoch. The small, medium, and large models were trained for 15 epochs with a batch size of 64,000 atoms, corresponding to an average of 1,454 systems per batch. The extra-small model was trained for 10 epochs. The medium model was adapted to the domain by adding one additional message-passing layer; otherwise, the model hyperparameters correspond to those used in the OMat training exercises.
The conservative models were fine-tuned for 7 epochs with a batch size of 128,000 atoms and the learning rate of 0.0002 was reduced to 0.0001 for the fine-tuning. We note that all models (and especially the conservatives models) would benefit from additional training, but we were constrained by available GPU resources and we plan to publish improved models in the future. We evaluated all models on the OMol validation splits without rotational symmetrization. More extensive metrics will be published on the official benchmark website, and the models (as well as newer versions) will be available as part of the upet repository.
[…]
(a) Conservative models
Model Energy MAE Energy MAE/atom Force MAE
PET-S 0.438 0.0040 0.0160
PET-M 0.153 0.0014 0.0078
PET-L 0.134 0.0013 0.0071
(b) Non-conservative models
Model Energy MAE Energy MAE/atom Force MAE
PET-XS 0.191 0.0024 0.0212
PET-S 0.166 0.0021 0.0197
PET-M 0.0798 0.0010 0.0086
PET-L 0.0707 0.00095 0.0072
Appendix H Details on geometry optimization and phonon calculations